OpenStudy (anonymous):

hog(x)=h(g(x)) h(x)=squared of x g(x)=4/(x-9)

OpenStudy (anonymous):

Replace the x in h(x) with g(x). \[h(g(x))=\left( \frac{4}{x-9}\right)^2=\frac{16}{(x-9)^2}\]